OpenStudy (anonymous):

The smallest angle in a triangle measures 90* less than the largest angle . The middle angle measures 60* less than the largest . Find the measure of each angle

OpenStudy (bahrom7893):

Okay let the smallest angle be x, middle angle be y and the largest one be z

OpenStudy (anonymous):

ok

OpenStudy (bahrom7893):

So: x = z-90 y - z - 60

OpenStudy (anonymous):

ok

OpenStudy (bahrom7893):

...thinking...

OpenStudy (anonymous):

now

OpenStudy (bahrom7893):

Oh and finally: x+y+z = 180

OpenStudy (bahrom7893):

plug in ur x and y values

OpenStudy (anonymous):

so is it 180-150?

OpenStudy (bahrom7893):

x = z-90 y = z - 60 x+y+z=180 z-90 + z-60 + z = 180

OpenStudy (bahrom7893):

3z - 150 = 180 3z = 180 + 150 3z = 330 z = 110

OpenStudy (bahrom7893):

x = z - 90 = 110 - 90 = 20 y = z - 60 = 110 - 60 = 50

OpenStudy (anonymous):

im so confused

OpenStudy (bahrom7893):

how? tell me where u are confused.

OpenStudy (anonymous):

well you wrote z=110 and then you put like 5 more steps. how long did it take you to learn this I seriously want to get this down.

OpenStudy (bahrom7893):

Okay here look: I said let x be the smallest, y be the middle, and z be the largest smallest is 90 less than the largest, therefore: x = z - 90 middle is 60 degrees less than the larges therefore: y = z-60

OpenStudy (bahrom7893):

Now sum of all angles in a triangle is given to be 180. That u have to know just memorize it.. sum is 180, so: smallest + medium + largest = 180 x + y + z = 180 Plug in what you had for x and y, remember how we let: x = z - 90, y = z - 60 (z - 90) + (z - 60) + z = 180

OpenStudy (bahrom7893):

Simplify that a bit further: z - 90 + z - 60 + z = 180 z + z + z - 60 - 90 = 180 3z - 150 = 180

OpenStudy (bahrom7893):

3z - 150 + 150 = 180 + 150 3z + 0 = 180 + 150 3z = 330 Divide everything by 3: z = 110

OpenStudy (bahrom7893):

Now you know the largest angle: z = 110 Smallest is 90 less than largest: x = z - 90 = 110 - 90 = 20 Medium is 60 less than largest: y = z - 60 = 110 - 60 = 50